Post adjusting entries, using the following information:

  1. Ryce paid cash for the following utility bills for the month: Light Company $13,000. The actual light bill received from Light Company is $10,000.
  2. Ryce stated business with a G Wagon valued at $1,200,000. Depreciation for the G Wagon has not been taken into account for the month. The business uses the reducing balance method to depreciate motor vehicles at a rate of 12 percent (12%) per annum.
  3. Ryce purchased store furniture from Royal Cabinets Limited using a cheque amounting to $120,000. Also, Ryce paid Telesales Electronics $15,000 with a cheque to set up her point-of-sale machine at the store. The furniture and point-of-sale machine in the business are to be depreciated using the straight-line method at a rate of 8%. Depreciation for furniture and point-of-sale the machine has not been taken into account.
1
Expert's answer
2021-06-27T17:30:01-0400

1.Debit settlements with suppliers credit cash register - 13 000

Debit main production credit settlements with suppliers - 10000

Debit cash register credit settlements with suppliers - 3000


2.It is necessary to add additional depreciation:

Debit main production of Credit depreciation - 12 000


"\frac{1 200 00\times0.12}{12}=12000"


3.t is necessary to add additional depreciation:

Debit main production of Credit depreciation - 960

Debit main production of Credit depreciation - 100


"\frac{15000\times0.08}{12}=100"



"\frac{120000\times0.08}{12}=960"